Please refer to the background information below to answer question 18 to 24. Utopia doctors immunize children against measles, a highly contagious and potentially deadly virus. The demand and supply curves for the immunization are given by the following equations. Demand: 200-P thousand immunizations per year Supply: Q = 8P - 45 thousand immunizations per year but the social marginal benefit (SMB) is given by the following equation Q=200 – 0.1 x SMB thousand immunizations per year and the private marginal cost and social marginal cost are identical. 18. If the market is left unregulated, the market equilibrium level of immunization is Answer18 ] thousand immunizations per year. 19. The total economic surplus to the society from the market equilibrium level of immunization is
1
Expert's answer
2020-12-21T10:57:22-0500

18. If the market is left unregulated, the market equilibrium level of immunization is:

Qd = Qs,

"200 - P = 8P - 45,"

P = 27.22,

Q = 200 - 27.22 = 172.88 thousand immunizations per year.

19. The total economic surplus to the society from the market equilibrium level of immunization is:

"TS = CS + PS = 0.5\u00d7(200 - 27.22)\u00d7172.88 + 0.5\u00d7(27.22 - 5.625)\u00d7172.88 = 16,801.78."